Optical Stark Spectroscopy of CaCH_{3}

A. J. Marr and T. C. Steimle, Dept. of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287-1604.

An intense pulsed supersonic molecular beam sample of CaCH_{3} was produced using a laser ablation/reaction source. The high resolution (50 MHz FWHM) laser induced fluorescence spectrum of the (K′= 1 - K′′= 0) Ã^{2}E-X̃^{2}A_{1} band system (origin = 14743 cm^{−1}) was recorded and fitted to produce an improved set of spectroscopic parameters. The resultant parameters were in agreement with those obtained in a previous Doppler limited study^{a}. A number of the transitions involving the lowest J levels were also recorded in the presence of variable static electric field. The Stark shifts and splittings were fitted to obtain dipole moments of 2.62(3) D and 1.69(1) D for the X̃^{2}A_{1} and Ã^{2}E states, respectively. The dipole moments were predicted using an electrostatic polarizability model^{b}^{,}^{c} that had successfully predicted the dipole moments of CaCCH ^{d}. The predicted values, 1.78 D X̃^{2}A_{1} and 1.20 D Ã^{2}E, are in reasonable agreement with experiment.

aC. R. Brazier and P. F. Bernath, J. Chem. Phys. 91, 4548 (1989).

bT. Törring, W. E. Ernst and J. Kländer, J. Chem. Phys. 90, 4927 (1989).

cJ. M. Mestdagh and J.-P. Visticot, Chem. Phys. 155, 19 (1991).

dA. J. Marr, J. Perry and T. C. Steimle, J. Chem. Phys. 103, 3861 (1995).